### v1

@@ 5.1.8 Miscellaneous
Open-Loop Optimal Controller | Wang et al. (2021b) | CRT (DER) | Reaching, Fetching | Numerical | ||||||||||||||||||||
Doroudchi et al. (2018) | EBRT | Shape Regulation | Numerical | |||||||||||||||||||||
Synergetic Control | Hashemi et al. (2023) | CRT | Tip Position TT | Numerical |
5.1.9 Shape/State Estimation
As seen in Sec. 5.1.1 and in Fig. 7, a general MB shape controller requires feedback in terms of configuration vector and its time derivative . Depending on the spatial discretization, the vector could be difficult or impossible to measure. To tackle this issue, a general MB control framework requires an Estimator, which maps the measures in the estimated state . For continuum and soft robots, this problem is referred to as Shape/State Estimation or Shape/State Reconstruction. Currently, most Shape Estimation methods use geometrical models or Computer Vision algorithms (e.g., (Bezawada et al., 2022; AlBeladi et al., 2021)). However, incorporating prior knowledge of rod-based models can significantly enhance performance.
@@ 5.1.9 Shape/State Estimation
In the pioneering work of Trivedi and Rahn (2014), the Authors suggest three approaches for Shape Estimation, based on CRT: (i) mounting load cells at the base, (ii) employing cable encoders, and (iii) mounting an inclinometer at the end of each piece. A maximum tip position error of is observed in simulations. Lastly, they use the OctArm VI (Grissom et al., 2006) to experimentally validate all three approaches. Another study proposes a real-time shape-estimation method based on the force-torque for CTR measured at the tubes’ basis (Donat et al., 2021). It extends a shape estimation algorithm for elastic Kirchhoff rods. They modeled a CTR combining planar PCC segments lying on different equilibrium planes. The approach is evaluated with single and two combined additively manufactured tubes at high frequency, achieving a mean deviation of mm along the tube. Differently, Lilge et al. (2022) develop a model for shape and strain estimation of a continuum cable-driven robot using a Gaussian Process regression, a mathematical tool for estimating continuous-time trajectories in . The idea involves substituting (i) time with arc-length and (ii) kinematic and dynamic laws based on CRT. This method efficiently estimates the robot’s shape using noisy measurements from sensors such as strain gauges, tracking coils, and an external camera. The real-world experiments provide excellent performances, with an average tip pose error of mm and . Finally, Yousefi et al. (2023) propose an optimization-based method that simultaneously estimates the shape and the forces acting on a continuum robot by employing the quasi-static CRT. Magnetic localization determines the position of multiple robot points. The method estimates the robot’s shape and force in a wide range of conditions validated experimentally, with and without knowledge of the contact positions.
5.2 Learning-based Controllers
@@
Learning-based controllers leverage the physics of rod models and utilize Machine Learning (ML) to derive parameterized control policies , which provide and end-to-end mapping between desired poses and sensor measurements to motor commands :
The vector of policy parameters are the weights of ANN trained via Supervised Learning (SL) or RL.
We report learning-based controllers in chronological order to highlight the evolution of the literature.
5.2.1 Supervised Learning
SL is viable to address the soft robotic control problem for simple tasks. It relies on pseudo-random motion data of actuations and corresponding task space to train an inverse model (50).
@@ 5.2.1 Supervised Learning
Initial learning strategies used rod models to rapidly prototype controllers. For instance, Thuruthel et al. (2017) developed dynamic models of CSM using recurrent neural networks and presented a trajectory optimization method for task space control. They validated the controller on a Cosserat rod with PCS parameterization and a pneumatic CSM. Recently, Wang and Rojas (2022) developed a closed-loop controller of continuum robots, employing RNN to approximate the forward and inverse dynamics. They extended the CRT in (Till et al., 2019) to account for spine compression and collect motion data for training their forward models. A non-parametric Gaussian process regression compensates for the error between the real robot and the RNN. A hybrid controller alternates between two stand-alone policies. They validated the control architecture on reaching and tracking tasks using a tendon-driven continuum robot with 3D-printed compressible spines. The results show that combining simulated and experimental data improves the control accuracy.
5.2.2 Reinforcement Learning
@@
RL is one of the main ML paradigms and a versatile framework suited for decision-making and control (Sutton and Barto, 2018). The two fundamental objects of an RL problem are the agent and the environment. In soft robotic control, the agent is a learning-based controller implementing an ANN, whereas the environment is everything outside the agent, comprising the soft robot and other objects. Fig. 8 depicts the agent-environment interaction highlighting the role of ANN and rod models. The agent and the environment interact at discrete time steps . At every interaction step, the agent observes the (possibly partial) state of the environment including sensor readings and task information. The agent acts according to a policy to select actions taking the form of low-level controls like torques or pressures. A policy can be a deterministic mapping from states to actions, . Alternatively, a policy can be stochastic, mapping a state to a distribution over actions . The environment transitions into another state according to the dynamics of the rod model. The agent also perceives a scalar reward signal from the environment, indicating the goodness of the current state specifying the control task. The goal of the agent is to maximize its return, the cumulative sum of rewards:
@@
The following deep RL contributions are summarized chronologically in Table 6. Satheeshbabu et al. (2019) applied deep Q-learning (DQL) with experience replay for the quasi-static position control of a pneumatic CSM capable of bending and twisting. An open-loop policy was trained in simulation utilizing a static Cosserat rod. They validate it in simulation and on the physical platform, subject to various external loads. To cope with the high-dimensional state-action space in soft robotics, DQL (Mnih et al., 2013) employs an ANN to approximate the value function of state-action pairs, defined as
A deterministic policy can leverage to select the best action in a state as .
@@
The following work increased the dexterity of the CSM and trained a closed-loop controller for precise quasi-static positioning via deep deterministic policy gradient (DDPG) (Satheeshbabu et al., 2020). Experiments in simulation investigated the robustness of the control policy for effects of loading, reachability, and workspace discontinuity. Finally, the controller was deployed on the prototype. The DDPG algorithm also derived a closed-loop controller for the quasi-static positioning of a pneumatic CSM modeled with the KRT (Uppalapati et al., 2020). Interestingly, the CSM was integrated into a mobile platform including a rigid arm and a sensorized gripper. The system was teleoperated and validated in the agricultural task of picking berries using different maneuvering strategies.
@@
The potential of continuum mechanical models is fully exploited in physical interaction tasks. Naughton et al. (2021) effectively applied several deep RL algorithms to train various control policies. The control tasks included point reaching, trajectory tracking, and maneuvering through structured and unstructured obstacles. The controllers were learned and tested in simulation using a synthetic CSM based on a dynamic CRT. The rod was actuated by applying distributed internal torques, modeled via splines characterized by control points and vanishing values at the rod ends. The work showcased a captivating interaction between the rod and the environment. In particular, the maneuvering task required the free end of the rod to reach a position behind fixed obstacles. The task was solved with a reward function that only considered the distance to the target and a penalty term for the occurrence of numerical errors
@@
Interestingly, the rod learned to navigate and interact with the obstacles without an explicit reward term. Finally, Alessi et al. (2023b) applied proximal policy optimization (PPO) to train a closed-loop position control policy for dynamic trajectory tracking with a pneumatic CSM. The controller was trained in simulation leveraging a dynamic Cosserat rod model of the CSM (Alessi et al., 2023a). After training, simulation tests evaluated the generalization capabilities of the policy to new observations, dynamics, and tasks. The experiments included tracking new trajectories subject to unknown external forces or using different material properties (e.g., ). In addition, the learned controller was transferred without retraining to intercept a moving object. The policy has shown good generalization capabilities for all four tests. Subsequently, Alessi et al. (2024) elaborated on the previous simulated environment to account for the dynamics of a dexterous CSM and its interaction with objects. They developed a learning-based control architecture for pose/force control, which enabled dynamic pushing with CSM in the real world. To mitigate the significant sim-to-real gap in soft robotics, they originally applied domain randomization (DR), an effective sim-to-real technique (Zhao et al., 2020b), emphasizing soft materials.
@@
Open loop ANN | Satheeshbabu et al. (2019) | CRT () | DQL | Tip Position TT (way-point) | (None) | ||||||
Closed loop ANN | Satheeshbabu et al. (2020) | CRT () | DDPG | Tip Position TT (way-point) | (None) | ||||||
Uppalapati et al. (2020) | KRT () | DDPG | Tip Position TT (way-point) | (None) | |||||||
@@
Naughton et al. (2021) | CRT () | Several | Tip Pose TT, Obstacle maneuvering | ||||||||
@@
Alessi et al. (2023b) | CRT () | PPO | Tip Position TT, Object interception | ||||||||
Alessi et al. (2024) | CRT () | PPO | Tip Pose/Force for Pushing | (DR) |
5.3 Discussion on Rod-based Controllers
5.3.1 On the Model-based Controllers

### v2

@@ 7.4 Robust and Optimal Control
| Open-Loop Optimal Controller | wang2021optimal | CRT (DER) | Reaching, Fetching | Numerical | ||||||||||||||||||||
| doroudchi2018decentralized | EBRT | Shape Regulation | Numerical | |||||||||||||||||||||
| Synergetic Control | hashemi2023robust | CRT | Tip Position TT | Numerical |
7.5 Shape/State Estimation
@@
A general MB shape controller requires feedback in terms of configuration vector and its time derivative (Figure 5). Depending on the spatial discretization, the vector could be difficult or impossible to measure. To tackle this issue, a general MB control framework requires an Estimator, which maps the measures to the estimated state . For continuum and soft robots, this problem is referred to as Shape/State Estimation or Shape/State Reconstruction. Shape Estimation methods can use geometrical models or Computer Vision algorithms (bezawada2022shape; albeladi2021vision). However, the following works show that incorporating prior knowledge of rod-based models can significantly enhance performance. The pioneering work of trivedi2014model suggests three approaches for Shape Estimation, based on CRT: (i) mounting load cells at the base, (ii) employing cable encoders, and (iii) mounting an inclinometer at the end of each piece. A maximum tip position error of is observed in simulations. All three approaches were then experimentally validated on the OctArm VI (grissom2006design). Another study by donat2021real proposes a real-time shape-estimation method based on the force-torque measured at the basis of the tubes for a CTR. It extends a shape estimation algorithm for elastic Kirchhoff rods. They modeled a CTR combining planar PCC segments lying on different equilibrium planes. The approach is evaluated with single and two combined additively manufactured tubes at high frequency, achieving a mean deviation of mm along the tube. Differently, lilge2022continuum developed a model for shape and strain estimation of a continuum cable-driven robot using a Gaussian Process regression to estimate continuous trajectories in . The idea involves substituting (i) time with arc-length and (ii) kinematic and dynamic laws based on CRT. This method efficiently estimates the robot’s shape using noisy measurements from sensors such as strain gauges, tracking coils, and an external camera. The real-world experiments provide excellent performances, with an average tip pose error of mm and . In parallel, aloi2022estimating developed a model-based framework for estimating distributed contact forces directly from shape measurements. The method involved fitting a Cosserat rod kinetostatic model with Gaussian load parameterization and solving it via an Extended Kalman Filter, enabling real-time, sensorless force estimation with sub-Newton accuracy. State estimation under uncertainty was also addressed in reconfigurable parallel continuum systems (anderson2017continuum), applying Cosserat rod modeling, multi-point constraints, and advanced filtering to improve tip accuracy and enable load inference. ferguson2024unified presents a surgical continuum manipulator with integrated FBG sensing for real-time 3D shape reconstruction, enabling sub-millimeter navigation without radiation in complex surgical settings. shi2016shape provided a comprehensive review of shape-sensing modalities for surgical continuum robots, highlighting complementary strengths of electromagnetic, optical, FBG, and vision-based methods. On the control side, underactuated discrete rod models were leveraged to achieve precise task-space trajectory tracking through FL and robust sliding-mode strategies, all while relying on minimal sensing (rucker2022task). In addition, yousefi2023model propose an optimization-based method that simultaneously estimates the shape and the forces acting on a continuum robot by employing the quasi-static CRT. Magnetic localization determines the position of multiple robot points. The method estimates the robot’s shape and force in a wide range of conditions validated experimentally, with and without knowledge of the contact positions. Recently, zheng2024full; zheng2024estimating proposed the Cosserat theoretic boundary observer, a mechanics-based dynamic state estimation algorithm to recover the infinite-dimensional robot states by measuring the tip velocity twist and the tip pose. Finally, feliu2024dynamic presents a novel method for estimating the dynamic state (pose and velocity) of tendon-actuated CSM. The approach uniquely combines a GVS Cosserat rod model with a nonlinear State-Dependent Riccati Equation observer. A key advantage of this technique is its ability to estimate the state using only internal sensor data, such as tendon displacements and actuator forces. The method was validated both in simulation and on a prototype, achieving an average dynamic tip position error of 1.79 cm, which is about 7%L.
8 Learning-based Controllers for Rod Models
@@
ML methods are becoming a trend to control CSM (george2018control; wang2021survey; wang2022control; falotico2024learning). Learning-based controllers leverage the physics of rod models and utilize ML to derive parameterized control policies , which provide and end-to-end mapping between desired poses and sensor measurements to motor commands :
The vector of policy parameters represents the weights of ANN, which can be trained with Supervised Learning (SL) or RL. We report learning-based controllers in chronological order to highlight the evolution of this growing field.
8.1 Supervised Learning
@@
SL in simulation is a viable approach to address simple tasks with soft robots. It relies on pseudo-random motion data of actuations and corresponding task space to train an inverse model (47). For instance, thuruthel2017learning developed dynamic models of CSM using recurrent neural networks and presented a trajectory optimization method for task space control. They validated the controller on a Cosserat rod with PCS parameterization and a pneumatic CSM. Moreover, wang2022data developed a closed-loop controller of continuum robots, employing RNN to approximate the forward and inverse dynamics. They extended the model in (till2019real) to account for spine compression and collect motion data for training their forward models. A non-parametric Gaussian process regression compensates for the error between the real robot and the RNN. Then, a hybrid controller alternates between two stand-alone policies. They validated the control architecture on reaching and tracking tasks using a tendon-driven continuum robot. Control accuracy improved by combining simulated and experimental data. Recently, generative models were employed for the first time to achieve domain translation on a soft robot, supported by a Cosserat rod model (kushawaha2025domain). However, this was validated only in simulation for simple tracking tasks and required a substantial computational effort.
8.2 Reinforcement Learning
@@
RL is one of the main ML paradigms and a versatile framework suited for decision-making and control (sutton2018reinforcement). The two fundamental objects of an RL problem are the agent and the environment. In soft robotic control, the agent is a learning-based controller implementing an ANN, whereas the environment is everything outside the agent, comprising the soft robot and other objects. Figure 6 depicts the agent-environment interaction highlighting the role of ANN and rod models. The agent and the environment interact at discrete time steps . At every interaction step, the agent observes the (possibly partial) state of the environment including sensor readings and task information. The agent acts according to a policy to select actions taking the form of low-level controls like torques or pressures. A policy can be a deterministic mapping from states to actions, . Alternatively, a policy can be stochastic, mapping a state to a distribution over actions . The environment transitions into another state according to the dynamics of the rod model. The agent also perceives a scalar reward signal from the environment, indicating the goodness of the current state specifying the control task. The goal of the agent is to maximize its return, the cumulative sum of rewards:
@@
We review the following deep RL contributions in chronological order. These are also summarized in Table 4. satheeshbabu2019open applied deep Q-learning (DQL) with experience replay for the quasi-static position control of a pneumatic CSM that can bend and twist. The policy was open-loop and trained on a simulation of a static Cosserat rod. They validate it in simulation and on the physical platform, subject to various external loads. DQL can cope with the challenging state-action space of soft robots by employing an ANN to approximate the value function of state-action pairs (mnih2013playing), which is defined as
@@
Then, a deterministic policy can leverage to select the best action in a state as . The work was extended by increasing the dexterity of the CSM and developing a closed-loop policy, still for precise quasi-static positioning, via deep deterministic policy gradient (DDPG) (satheeshbabu2020continuous). They investigated the robustness of the control policy to loading and workspace discontinuity, and deployed the policy on hardware. The DDPG algorithm also derived a closed-loop controller for the quasi-static positioning of a pneumatic CSM modeled with the KRT (uppalapati2020berry). Interestingly, the CSM was integrated into a mobile platform including a rigid arm and a sensorized gripper. The system was teleoperated and validated in the agricultural task of picking berries using different maneuvering strategies.
@@
The potential of continuum mechanical models is fully exploited in physical interaction tasks. naughton2021elastica effectively applied several deep RL algorithms to train various control policies. They explored point reaching, trajectory tracking, and maneuvering through structured and unstructured obstacles. The controllers were learned and tested in simulation using a synthetic CSM based on a dynamic CRT. The rod was actuated by applying distributed internal torques, modeled via splines characterized by control points and vanishing values at the rod ends. The work showcased a captivating interaction between the rod and the environment. In particular, the maneuvering task required the free end of the rod to reach a position behind fixed obstacles. The task was solved with a reward function that only considered the distance to the target and a penalty term for the occurrence of numerical errors
@@
Interestingly, the rod learned to navigate and interact with the obstacles without an explicit reward term. The generalization capabilities of learned controllers are vital for successful deployment in the real world. In this regard, alessi2023learning applied proximal policy optimization (PPO) to train a closed-loop position control policy for dynamic trajectory tracking with a pneumatic CSM. The controller was trained leveraging a dynamic Cosserat rod model of the soft robot (alessi2023ablation). After training, four simulation tests evaluated how the policy generalized to new observations, dynamics, and tasks. The experiments included tracking new trajectories subject to unknown external forces or using different material properties (e.g., ). The policy generalized well within some boundaries and also transferred without retraining to intercept a moving object. Subsequently, alessi2024pushing extended the simulated environment to account for the dynamics of a dexterous CSM and its interaction with objects. The proposed closed-loop control architecture for pose/force control enabled dynamic pushing with CSM in the real world. They mitigated the significant sim-to-real gap in soft robotics with a novel adaptation of domain randomization (DR) (zhao2020sim) that emphasizes the randomization of soft material properties. This was the first sim-to-real transfer of a manipulation policy based on rod models.
@@
| Deep Q-Learning | satheeshbabu2019open | CRT (✗) | Tip Position TT (way-point) | ✓(None) | ||||
| Deep Deterministic Policy Gradient | satheeshbabu2020continuous | CRT (✗) | Tip Position TT (way-point) | ✓(None) | ||||
| uppalapati2020berry | KRT (✗) | Tip Position TT (way-point) | ✓(None) | |||||
@@
| Trust Region Policy Optimization | naughton2021elastica | CRT (✓) | Tip Pose TT, Obstacle maneuvering | ✗ | ||||
@@
| Proximal Policy Optimization | alessi2023learning | CRT (✓) | Tip Position TT, Object interception | ✗ | ||||
| alessi2024pushing | CRT (✓) | Pushing (Tip Pose/Force) | ✓(DR) |
9 Discussion and Emerging Challenges
9.1 On the Rod Theories
Following our mathematical treatment and review of rod models, we briefly reevaluate the strengths and limitations of the various rod theories within soft robotics. The Euler-Bernoulli model is simple, computationally light, and effective for slender, stiff structures subject to small bending deflections. Its nonlinear version makes it suitable for large deformations. However, it remains a planar model and neglects shear as the cross-sections remain perpendicular to the centerline. This assumption makes it unsuitable for modeling interactions with the environment in soft robotics. Nonetheless, some works extended its applicability beyond bending by coupling it with other models. The Timoshenko model improves on this by including shear deformation. However, Timoshenko is less suited to nonlinear actuator paths that induce torsion. To this end, the Kirchhoff–Love theory handles large rotations and displacements but assumes the rod is inextensible and unshearable. This assumption limits its ability to model extensible pneumatic actuators. Finally, the Cosserat-Reissner model and its extensions are the most general and suited for continuum soft robots. They can handle large deformations, shear, torsion, and stretching. However, they require more computational effort. Overall, each rod theory has its strengths and weaknesses. The choice depends on the system’s behavior and the level of detail needed. We suggest opting for a Cosserat rod model and tuning the space and time discretization to match the desired tradeoff between efficiency and accuracy.
9.2 On the Deformations
Researchers developed numerous rod-based models reformulating and adapting classical rod theories in the soft robotic field. Herein, we reviewed diverse models, which described different combinations of strain modes, giving rise to nine deformation classes.
The most common class in literature is Bend, with about a third of the examined papers. This popularity is due to the propensity of slender bodies to bend. In fact, due to the assumption of , with cross-section diameter, the bending stiffness turns out smaller than the stretching or shear ones, easing the bending deformation. In addition, most used actuator paths are linear, which excite predominantly bending modes. These reasons justify the approximation of considering only the bending mode. Furthermore, we can observe that most models in the Bend class leverage the EBRT. However, as summarized in Table 2, several models employed the general CRT. While this represents a computational surplus for trivial bending structures, the generalization of Cosserat models pays off if the model gets deployed in physical interaction tasks.
The next class is All Deformation, with about a quarter of the examined papers. As expected, only models based on the CRT achieve all deformations (Table 2). This group of models received a deep study in the last decade, thanks to the rising interest in physical interaction with unstructured environments where all strain modes are significant. Indeed, these models are useful to describe accurately the deformations of slender robots caused by friction (e.g., anisotropic friction for locomotors) or by different mediums (e.g., water). In addition, the most common actuation sources (i.e., pneumatic and cables) excite elongation and compression, causing a significant change in the dimension workspace.
The third class accounting for about a fifth of the investigated studies is Bend & Twist. These models mainly describe robots in which the unshearability and the unstretchability approximations hold (Table 2). Including the twisting deformation allows the representation of backbones with generic 3D curves. However, the twisting deformation can be excited only by specific actuator paths (e.g., helicoidal cables) or asymmetric actuators.
Another significant class is Bend & Stretch thanks to the wide diffusion of linear extensible soft pneumatic actuators and tendon-driven robots. Such actuators extend or compress upon solicitation, which induces the continuum body to stretch and bend thanks to their radial disposition on the cross-section.
The remaining five categories are less explored. While the Bend, Twist & Stretch is a particular case of the All Deformations class, the other classes comprehend fewer works. The lack of work is due to the difficulty of exciting these specific strain modes with the current actuation technologies or the small number of applications in which the assumptions are valid. In addition, fewer studies that explored shearing per se could be due to the tendency to consider shearing an undesirable or negligible deformation. However, shearing could be significant when there is a prominent environment interaction, as observed in peristaltic locomotion (hemingway2021continuous) or complex manipulation (grazioso2019geometrically). Shearing also becomes important when the aspect ratio of the CSM becomes less slender and more tick.
9.3 On the Spatial Discretization
A characterizing feature of rod models is the spatial discretization method, which significantly impacts the model’s accuracy and complexity. For simulation and control purposes, the discrete model must be computationally efficient and as accurate as possible. In the analyzed literature, two primary discretization approaches are commonly employed: Discrete Elastic Rod (DER) (Sec. 4.2.4) and Strain Parameterization (Sec. 4.2.3).
The DER method discretizes the rod’s length into nodes connected by rigid links, representing the rod CSM as a set of discrete poses. The accuracy of the solution relative to the continuous rod depends directly on the number of nodes. In contrast, Strain Parameterization discretizes the rod in the configuration space, constraining the robot’s backbone to a subset of admissible spatial curves. The latter approach permits the neglect of specific strain modes without imposing additional constraints on the EoM, as required in the DER formulation, thereby enhancing computational efficiency. Furthermore, Strain Parameterization naturally allows the forward dynamics to be expressed in Lagrangian form, which is particularly advantageous for control design. Both approaches are widely used and integrated into dedicated simulators, such as SoRoSim (mathew2022sorosim) for Strain Parameterization and PyElastica (naughton2021elastica) for DER.
Straddling geometric and mechanical models, some examined works proposed to fit discrete points with functions, such as PH or Euler’s spirals, which is useful, especially for Shape Estimation.
9.4 On the Identification and Validation
to assess the correctness of the rod model,
to estimate the unknown parameters.
In the absence of precise FEM models, lumped-mass models with a high number of DoFs were employed. Otherwise, researchers resort to model optimization to match experimental data. In particular, when the robot morphology is not uniform, researchers usually estimate geometrical properties (e.g., cross-sectional area ) and material characteristics (e.g., elastic modulus , damping coefficient ). Experiments typically conducted are with the robot at different base orientations subject to external loads or at the tip. Most rod models are identified and validated through motion data collected with predefined actuation profiles . However, the literature needs rigorous mechanical tests to yield more accurate models, especially for unstructured physical interactions with the environment. Indeed, despite a careful model calibration on motion data, alessi2023ablation; alessi2024pushing observed real-world buckling during a manipulation task that was not present in simulation. Concerning performance metrics, rod models predicted experimental tip positions with Euclidean errors in the range . However, only a few studies assessed the angular errors, which play a role in dexterous manipulators. In summary, the wide use of experimental data suggests an increasing need to reduce the reality gap to develop more accurate controllers.
9.5 On the Impact of Materials and Manufacturing
Manufacturing techniques provided continuum and soft robots with unprecedented dexterity (wallin20183d). However, manufacturing uncertainties may impact model accuracy. Recent rod models addressed manufacturing uncertainties considering variations in material (alessi2023ablation) and geometric properties (eugster2022soft).
Research gaps persist despite these advancements. First, capturing hysteresis—a prevalent phenomenon in soft materials that affects the accuracy and predictability of the robot’s behavior—remains largely underexplored in rod models. Second, incorporating self-healing materials can potentially enhance the durability and longevity of soft robots (terryn2017self). However, integrating the dynamic restoring process into rod models should still be investigated. Filling these gaps could advance the capabilities of continuum and soft robots, enabling their integration into diverse applications and real-world tasks.
9.6 On the Model-based Controllers